OpenStudy (anonymous):

the equation of the line with slope -2 that goes through (-3,-7)

OpenStudy (anonymous):

y=mx+b y=-2-7

OpenStudy (anonymous):

y=-2x-7 m=slope, b=y-intercept

OpenStudy (anonymous):

Thank u